OpenStudy (anonymous):

Explain why tan 45degrees= 1?

OpenStudy (anonymous):

tan(x) = sin(x)/ cos(x)

OpenStudy (anonymous):

\[\tan(\theta)=\frac{ \sin(\theta) }{ \cos(\theta) } \implies \tan(45)=\frac{ \sin(45) }{ \cos(45) }=\frac{ \sqrt2 /2 }{ \sqrt2 / 2 }=1\] @Maquela